### v2

Abstract
Third-Party Risk Assessment (TPRA) is a core cybersecurity practice for evaluating suppliers against standards such as ISO/IEC 27001 and NIST. TPRA questionnaires are typically drawn from large repositories of security and compliance questions, yet tailoring assessments to organizational needs remains a largely manual process. Existing retrieval approaches rely on keyword or surface-level similarity, which often fails to capture implicit assessment scope and control semantics.
This paper explores strategies for organizing and retrieving TPRA cybersecurity questions using semantic labels that describe both control domains and assessment scope. We compare direct question-level labeling with a Large Language Model (LLM) against a hybrid semi-supervised semantic labeling (SSSL) pipeline that clusters questions in embedding space, labels a small representative subset using an LLM, and propagates labels to remaining questions using k-Nearest Neighbors; we also compare downstream retrieval based on direct question similarity versus retrieval in the label space. We find that semantic labels can improve retrieval alignment when labels are discriminative and consistent, and that SSSL can generalize labels from a small labeled subset to large repositories while substantially reducing LLM usage and cost.
1 Introduction
As organizations increasingly rely on third-party providers such as cloud platforms and SaaS vendors, Third-Party Risk Assessment (TPRA) has become a core component of cybersecurity risk management and regulatory compliance. TPRA typically relies on structured questionnaires derived from standards such as ISO/IEC 27001 and NIST [8, 15]. Organizations select relevant questions from internal repositories based on a supplier’s scope and risk profile, evaluate the supplier’s responses, and decide whether to proceed with onboarding or request remediation [11, 16].
While significant effort is devoted to analyzing supplier responses, a major operational bottleneck arises earlier in the process: selecting relevant questions. Large organizations maintain extensive repositories of compliance questions accumulated across standards and assessment campaigns, yet these repositories are typically unstructured and curated manually [14]. As a result, question selection is repetitive, time-consuming, and difficult to scale.
Most automated approaches frame question selection as a retrieval problem and rely on text similarity [23], including cosine similarity over dense embeddings [21]. However, such methods operate purely on semantic similarity and lack explicit knowledge of the underlying control domain (e.g., access control, incident response) or assessment scope (e.g., existence verification, enforcement on critical systems). When these aspects are implicit, retrieval often yields overly generic questions that do not align with the assessment scope.
An alternative is to enrich compliance repositories with semantic labels capturing both control domain and scope, enabling label-based retrieval. However, most repositories lack such labels, and there is no established methodology for deriving them automatically. Existing work has explored automated labeling using topic modeling [9] and clustering-based methods [5]. More recently, Large Language Models (LLMs) have shown strong performance as annotators and can generate human-readable labels without predefined taxonomies [24, 22]. Nevertheless, using LLMs as the primary labeling mechanism for large and evolving repositories is costly and sensitive to prompt variability.
To explore this design space, we propose a hybrid Semi-Supervised Semantic Labeling (SSSL) framework for TPRA compliance questions. We first cluster questions using embedding-based similarity to provide the LLM with richer context and induce more broadly applicable labels. We then label a small set of representative questions with LLM assistance to form an interpretable seed label set, and propagate these labels to the remaining questions using a k-Nearest Neighbors (kNN) classifier, minimizing repeated LLM inference. We evaluate SSSL against an LLM-only labeling baseline on label quality, generalization, and cost, and show up to 40% lower labeling cost and 33% reduction in runtime throughput while maintaining comparable generalization. Finally, we use the induced labels for repository-scale retrieval and find that label-based retrieval can better align selected questions with intended control domains and assessment scope than semantic similarity alone, although precision remains challenging in some cases.
2 Related Work
Recent work on Third-Party Risk Assessment (TPRA) automation spans both industrial systems and academic research. In practice, the main challenge is not the lack of assessment content, but the ability to organize, reuse, and adapt large question repositories. We group related work into three areas: TPRA systems, similarity-based retrieval methods, and text labeling approaches.
2.1 Third-Party Risk Assessment Systems
TPRA is commonly implemented within third-party risk management (TPRM) workflows that standardize vendor onboarding, questionnaire-based assessments, remediation tracking, and continuous monitoring [11]. Commercial platforms support these workflows through curated question libraries, predefined mappings to compliance frameworks, and reporting pipelines.
While effective for governance and auditability, these systems rely heavily on manually curated content and static configurations. Question repositories are typically accumulated across assessment campaigns, resulting in redundancy, inconsistent phrasing, and limited reuse. Adapting questionnaires to organization-specific needs or evolving risk contexts therefore requires significant manual effort, motivating automated yet interpretable approaches to question discovery and selection.
2.2 Similarity-Based Methods
Similarity-based retrieval ranks questions or controls by textual relevance. Traditional lexical methods depend on vocabulary overlap and are sensitive to phrasing variation [19]. Neural retrieval methods embed queries and documents into dense vector spaces, enabling semantic similarity matching [18, 10, 12].
Although effective for finding related questions, these methods do not explicitly model assessment intent, control domain, or evaluation scope. In practice, this often yields results that are topically related but misaligned with the intended assessment objective, limiting their usefulness for constructing structured TPRA questionnaires.
2.3 Text Labeling Methods
An alternative to direct retrieval is to enrich repositories with semantic labels. Unsupervised approaches, such as topic modeling and embedding-based clustering, attempt to discover latent structure without annotation. Classical topic models struggle with short compliance texts [2], while embedding-based clustering typically assigns each item to a single cluster, limiting representation of overlapping compliance concepts [5].
Weak supervision generates labels using heuristics or labeling functions [17], but requires a predefined label space and substantial upfront design effort. More recently, Large Language Models (LLMs) have been used to generate human-readable labels without fixed taxonomies [22]. However, LLM-based labeling is costly, sensitive to prompt design, and variable across runs [7]. Hybrid approaches reduce some of these limitations but often still rely on repeated LLM inference, constraining scalability in large or evolving repositories.
3 Hybrid Possibilistic Clustering with LLM-Assisted Labeling
This section delineates the SSSL framework. Section 3.1 introduces an overview of SSSL. Section 3.2 presents the first phase, which uses an LLM to discover semantic labels. Section 3.3 then describes how the generated labels are used to infer labels for new questions. Finally, Section 3.4 explains how these labels are leveraged for question retrieval.
3.1 SSSL Pipeline
We propose SSSL, a semi-supervised semantic labeling framework for TPRA compliance questions, which is illustrated in Figure 1. The first step in SSSL is repository construction, which transforms unlabeled data into labeled data. Specifically, we take as input unlabeled TPRA compliance questions and group them using semantic embeddings with overlapping cluster membership. An LLM then assigns cluster-level semantic labels, which are aggregated into multi-label annotations for each question (annotation phase). The resulting labeled repository enables k-NN–based label prediction for new questions without requiring repeated LLM inference (prediction phase).
The second step is query-time inference, which supports retrieval in the label space aligned with the supplier context in the user request, thereby enabling the retrieval of relevant TPRA questions (label-based retrieval phase).
Figure 1: Proposed SSSL pipeline
3.2 Annotation Phase
Let $Q=\{q_{i}\}_{i=1}^{n}$ be the input set of cybersecurity compliance assessment questions. Our goal is to assign each question a (possibly multi-label) set $L_{i}\subseteq L$, where $L=\{l_{j}\}_{j=1}^{m}$ is a free-form label inventory discovered by an LLM at the cluster level. We first cluster questions into overlapping groups $\{C_{z}\}_{z=1}^{s}$ and obtain labels $L^{\prime}_{z}$ for each group via a single LLM call. Each question then inherits the union of its groups’ labels: $L_{i}=\bigcup_{z:q_{i}\in C_{z}}L^{\prime}_{z}$.
Question Embeddings.
We represent each compliance question as a dense semantic vector using a pretrained sentence embedding model. Formally, for each $q_{i}\in Q$, we compute
$\mathbf{e}_{i}=f(q_{i}),\qquad\mathbf{e}_{i}\in\mathbb{R}^{d},$
where $f(\cdot)$ is fixed across all experiments [^1] We use the text-embedding-3-large, from openai https://openai.com/, embedding model because it provides one of the strongest embedding representations currently available, with an output vector dimension of 3072.. These embeddings enable similarity-based grouping while being robust to lexical variation common in compliance phrasing.
Possibilistic clustering.
To form semantically coherent groups without forcing exclusive assignments, we apply Possibilistic C-Means (PCM) [1, 13]. Unlike hard clustering, PCM assigns each question an independent membership score for each cluster, allowing overlap between groups. Let $k$ denote the number of clusters (we set k=|Q|1k=|Q|-1). PCM returns a possibility membership matrix
$\mathcal{P}\in\mathbb{R}^{k\times n},\qquad P_{c,i}\in[0,1],$
where $P_{c,i}$ indicates the degree to which question $q_{i}$ belongs to cluster $C_{c}$. [^2] PCM memberships are not normalized across clusters; we therefore treat them as degrees of association rather than probabilities.
To convert soft memberships into discrete (possibly overlapping) clusters, we threshold memberships on a per-cluster basis. Specifically, for each cluster $c$, we select a membership threshold $\tau_{c}$ automatically using elbow (knee) detection on the sorted membership scores $\{P_{c,i}\}_{i=1}^{n}$ [20]. This avoids dataset-specific manual tuning and adapts to the skewed similarity distributions often observed in compliance question repositories. When no clear elbow is detected, we conservatively fall back to the minimum membership value to avoid discarding potentially relevant associations.
Algorithm 1 summarizes this procedure and yields a set of overlapping clusters $\mathcal{C}=\{\mathcal{C}_{1},\ldots,\mathcal{C}_{k}\}$, where each $\mathcal{C}_{c}\subseteq Q$ is constructed independently. These clusters are then passed to the LLM-based annotator for label discovery.
Algorithm 1 Elbow-Based Possibilistic Question Grouping
1:Question $Q=\{q_{1},q_{2},\dots,q_{n}\}$, membership matrix $\mathcal{P}\in\mathbb{R}^{(k=|Q|-1)\times n}$
2:clusters $\mathcal{C}$
[…]
4:for each $i\in\{1,\dots,k\}$ do
[…]
7:if $k_{i}=\varnothing$ then
8:  $\tau_{i}\leftarrow\min(\mathbf{v}_{i})$
9:else
10:  $\tau_{i}\leftarrow\mathbf{v}_{i}[k_{i}]$
[…]
13:Return $\mathcal{C}$
LLM-assisted Labeling.
LLMs can produce high-quality, human-readable labels, but invoking them repeatedly at the question level is costly and may yield inconsistent label names across runs. We therefore restrict LLM usage to the cluster level: given the overlapping clusters produced by Algorithm 1, we invoke the LLM once per cluster to extract a small set of reusable semantic labels that are shared across the questions in that cluster. This design amortizes LLM cost over many questions and provides the model with additional context, which empirically improves label coherence compared to labeling questions independently.
To improve reproducibility and reduce output variance, we use a structured prompt template (Figure 2) [22, 3]. The template: (i) assigns a domain role (cybersecurity audit classification), (ii) constrains the task to cluster-common labels rather than question-specific details, (iii) enforces comparison across questions to identify shared themes, and (iv) restricts outputs to short noun-phrase labels with no explanations or predefined taxonomies. For each cluster $\mathcal{C}_{i}$, the LLM returns a label set $L_{i}$.
Figure 2: Structured prompt template used for cluster-level label extraction.
Because labels are produced at the cluster level, we derive per-question label sets by aggregating labels from all clusters containing the question. Concretely, each question $q$ inherits the union of labels from its assigned clusters:
$\mathcal{L}[q]=\bigcup_{i:q\in\mathcal{C}_{i}}L_{i}.$
Algorithm 2 implements this aggregation and yields a labeled repository $\mathcal{L}$ that is used in the prediction phase.
Algorithm 2 Collecting Labels per Question
1:Questions $Q=\{q_{1},\ldots,q_{n}\}$, clusters $\{\mathcal{C}_{i}\}_{i=1}^{k}$, cluster labels $\{L_{i}\}_{i=1}^{k}$
2:Per-question label mapping $\mathcal{L}:Q\rightarrow 2^{L}$
3:Initialize $\mathcal{L}[q]\leftarrow\emptyset$ for all $q\in Q$
4:for each $i\in\{1,\dots,k\}$ do
5:for each $q\in\mathcal{C}_{i}$ do
6:  $\mathcal{L}[q]\leftarrow\mathcal{L}[q]\cup L_{i}$
7:Return $\mathcal{L}$
3.3 Prediction Phase
The prediction phase assigns labels to a new compliance question using the labeled repository produced in the annotation phase, avoiding LLM inference under normal operation. Because the labeled set is typically small and evolves over time, we use instance-based learning with k-Nearest Neighbors (kNN), a non-parametric method that supports incremental updates without retraining [25]. Given a new question $q^{\prime}$, we compute its embedding $\mathbf{e}^{\prime}=f(q^{\prime})$ and retrieve the $k$ most similar labeled questions in the embedding space using cosine similarity.
To infer labels from the retrieved neighbors, we use a simple voting scheme. Each neighbor contributes its label set, and each label receives one vote per occurrence among the neighbors [4]. The predicted label set consists of the label(s) with the maximum vote count, allowing ties. To improve robustness, we suppress labels that appear in only one neighbor. If no label receives at least two votes, we treat $q^{\prime}$ as out-of-distribution and defer labeling to the LLM.
Algorithm 3 summarizes the kNN-based label propagation procedure used to predict labels for a new question.
Algorithm 3 kNN Label Selection for a New Question
1:Labeled questions $Q=\{q_{1},\ldots,q_{n}\}$, per-question labels $\mathcal{L}[q_{i}]$, embedding model $f(\cdot)$, number of neighbors $k$
2:New question $q^{\prime}$
3:Predicted label set $L(q^{\prime})$
[…]
5:for all $i\in\{1,\ldots,n\}$ do
[…]
8:Initialize vote map $S[\ell]\leftarrow 0$ for all labels $\ell$
9:for all $i\in\mathcal{N}$ do
10:for all $\ell\in\mathcal{L}[q_{i}]$ do
11:  $S[\ell]\leftarrow S[\ell]+1$
[…]
13:if $S_{\max}<2$ then
14:return $\textsc{LLMLabel}(q^{\prime})$ {OOD fallback}
[…]
16:return $L(q^{\prime})$
3.4 Label-based Retrieval Phase
Once questions are enriched with semantic labels, we perform retrieval in the label space rather than directly in the question-text space. Direct question-embedding similarity often returns items that are lexically or topically close but differ in assessment intent (e.g., policy requirements vs. operational evidence requests). Our labels make control domains and scope explicit, enabling more targeted question selection.
Let $\mathcal{R}=\{(q_{i},\mathcal{L}[q_{i}])\}_{i=1}^{n}$ be the labeled repository, where $\mathcal{L}[q_{i}]\subseteq L$ is the (multi-)label set assigned to question $q_{i}$, and each label $\ell\in L$ is a short natural-language descriptor (e.g., Access control, Incident response). We embed every label once using a sentence encoder $f(\cdot)$, yielding a label embedding $\mathbf{e}_{\ell}=f(\ell)$. Given a user request $u$, we compute its embedding $\mathbf{e}_{u}=f(u)$ and score each question by matching $\mathbf{e}_{u}$ against the embeddings of its associated labels:
$\mathrm{score}(u,q_{i})=\operatorname{Agg}_{\ell\in\mathcal{L}[q_{i}]}\cos(\mathbf{e}_{u},\mathbf{e}_{\ell}),$
where $\operatorname{Agg}$ aggregates label-level similarities (we use the mean; max is a strong alternative). We return the top-$r$ questions ranked by $\mathrm{score}(u,q_{i})$.
4 Evaluation
We evaluate whether the proposed SSSL framework reduces reliance on repeated LLM-based labeling for compliance self-assessment questionnaires, while enabling retrieval of questions based on reusable semantic labels rather than question-level similarity. To support reproducibility, we release an open-source implementation for research and non-commercial use, along with all datasets, evaluation results, and scripts at: https://github.com/NourEldin-Ali/semi-supervised-semantic-labeling.
4.1 Experimental Setup
Datasets. Our evaluation focuses on cybersecurity compliance self-assessment questionnaires, a common artifact in industry governance, risk, and compliance workflows. Public datasets of such questionnaires are scarce due to licensing restrictions; we therefore include the publicly available Consensus Assessments Initiative Questionnaire (CAIQ) published by the Cloud Security Alliance as a real-world benchmark. [^3] https://d1.awsstatic.com/whitepapers/compliance/CSA_Consensus_Assessments_Initiative_Questionnaire.pdf
To evaluate scalability across standards, we additionally construct synthetic questionnaires guided by publicly available control descriptions and industry mappings from widely adopted security frameworks (e.g., ISO/IEC 27001 family, ISO 27002, etc). Each dataset contains 150–250 short natural-language questions describing security, governance, or operational controls. Although synthetic, these datasets preserve two key properties observed in real compliance assessments: high semantic overlap across standards and the absence of a fixed labeling taxonomy.
Baselines and Metrics. We compare SSSL against (i) an LLM-based single-question annotator that assigns labels independently, and (ii) a KNN-based predictive labeling method using embedding similarity. All experiments use GPT-5.1.
In addition to label quality, we measure label generalizability, execution time, token consumption, and energy usage to capture the operational cost of each approach. [^4] Energy usage is measured using CodeCarbon: https://codecarbon.io/.
Evaluation Tasks. We evaluate performance along two practical dimensions.
Label quality evaluation assesses whether generated labels accurately capture question intent and remain reusable across semantically related questions, measuring label correctness, label generalization (abstraction beyond individual instances), and consistency (uniform and interpretable naming). Label quality is assessed using an LLM-as-a-Judge [6] with explicit scoring criteria, enabling scalable comparison across datasets. In addition to the time consumed using different methods and the tokens spent, we also consider the energy consumed. This can help us take into account a multidimensional evaluation of label generation and their quality.
Question retrieval evaluation examines whether semantic labels improve retrieval of questions sharing the same underlying control scope, compared to retrieval based solely on question-level embedding similarity.
4.2 Results & Discussion
Experiments are conducted across multiple datasets, and we observe consistent relative trends across all settings. [^5] Due to space limitations, we report only the average results in this paper; the full results are available in our GitHub repository.
Semantic Label Quality.
We first evaluate the quality of the semantic labels produced by different labeling strategies. Label quality is assessed using an LLM-based judge, which assigns a score from 1 to 5, where 5 indicates that the label fully and accurately captures the shared semantic intent of the associated questions.
Method Correctness Generalization Consistency
LLM 4.8 4.3 4.8
SSSL (LLM phase) 3.5 4.3 4.8
SSSL (kNN phase) 1.8 3.4 4.7
Table 1: Semantic label quality comparison evaluated using an LLM-based judge. Higher is better.
Table 1 shows that LLM-only labeling achieves the strongest correctness (4.8) but, as a per-question strategy, it may yield less stable labels due to independent annotation decisions. In contrast, the cluster-level LLM phase of SSSL leverages shared context within semantically coherent groups, producing highly consistent labels (4.8 consistency) with comparable generalization (4.3). Finally, the kNN-based propagation phase maintains near-identical consistency (4.7) while reducing correctness (1.8) and generalization (3.4), indicating that propagation is efficient but can be sensitive to boundary cases where nearest-neighbor transfer is less precise.
Computational Efficiency and Scalability.
We next compare computational cost across methods. Table 2 summarizes the number of labels produced, energy usage, token consumption, and runtime.
Method Labels # Energy (kWh) Tokens Time (s)
LLM 628 0.003 57146 322
SSSL (LLM phase) 131 0.002 34527 214
SSSL (kNN phase) 18 0.000002 0 0.22
Table 2: Computational cost comparison. Lower is better except for Labels #.
[…]
Label Selection Quality.
We then evaluate downstream label selection: given a high-level assessment intent, the system retrieves and ranks candidate semantic labels that best match the intent. Performance is measured using a score from 0 to 100, where 100 indicates a perfect match between the retrieved label and the intended assessment. We evaluate three label-selection queries: Q1 (Backup Frequency & Retention), Q2 (Monitoring & Incident Response), and Q3 (Backup Frequency & Retention + Monitoring & Incident Response), a composite query that requires selecting labels spanning both domains. Selection quality is evaluated using the same LLM-based judge.
Method Q1 Q2 Q3 Overall
BM25 28 78 70 58
Semantic similarity 58 90 62 70
Labeling Semantic Similarity 65 88 72 75
Table 3: Label selection quality evaluated using an LLM-based judge. Higher is better.
As shown in Table 3, BM25 performs unevenly across queries, reflecting its reliance on lexical overlap: it is comparatively strong on Q2 (78) where operational terms are often explicit, but substantially weaker on Q1 (28) where equivalent requirements may be expressed with diverse phrasing. Embedding-based semantic similarity improves overall performance (70), but still drops on the composite Q3 (62), suggesting that unconstrained similarity can retrieve labels that are related to only one part of the combined intent. Incorporating semantic labels into the similarity signal yields the best overall performance (75) and improves both the single-domain queries (Q1: 65; Q2: 88) and the multi-domain composite (Q3: 72).
Discussion.
Taken together, these results demonstrate that decoupling semantic label discovery from large-scale label assignment enables an effective and scalable labeling pipeline. The cluster-level LLM phase of SSSL constructs a consistent semantic label space, while the kNN phase enables fast, zero-token deployment by transferring labels via embedding similarity.
The drop in label quality in the kNN phase (Table 1; correctness 1.8 vs. 3.5 in the SSSL LLM phase) is expected in our cross-standard setting. In particular, we propagate labels learned from one cybersecurity standard to questionnaires aligned with another standard that does not fully overlap in scope and terminology. This partial mismatch can lead to (i) semantic drift, where the nearest neighbors are topically related but not aligned with the exact compliance intent, reducing correctness; and (ii) label-space contraction, where only a subset of the discovered labels are meaningfully transferable, reducing the number of distinct labels used in the kNN phase (Table 2; 18 labels vs. 131 in the SSSL LLM phase). In other words, kNN propagation is most effective when the target questionnaire is drawn from the same (or highly overlapping) control taxonomy; when applied across standards with incomplete overlap, fewer labels match cleanly and correctness degrades for boundary cases. Despite this limitation, the kNN phase provides substantial practical value by enabling near-instant label assignment (0.22 s; 0 tokens), making it suitable for rapid deployment and iterative updates when approximate labeling is acceptable or when a small amount of targeted re-annotation can correct mismatched regions.
One promising direction to address this limitation is to explore grouping at the label level rather than the question level. In other words, instead of clustering questions to generate labels, we could first generate labels for individual clusters and then group those labels into higher-level categories. While this would not reduce the initial cost of LLM labeling, it could expand the effective label space used during kNN-based propagation, potentially improving both correctness and label diversity. In particular, using grouped or generalized labels during propagation may enable more robust transfer across heterogeneous standards and reduce errors caused by overly narrow or underrepresented labels.
4.3 Threats to Validity
Our evaluation focuses on cybersecurity compliance questionnaires used in TPRA workflows. While these datasets reflect real industrial use, results may differ for other assessment or governance scenarios with distinct terminology, granularity, or structure. Evaluating SSSL on additional industrial domains is left for future work.
Our experiments rely on a single LLM for both semantic label discovery and LLM-based evaluation. Different LLMs may vary in labeling behavior and judgment criteria. A broader evaluation using multiple LLMs, both as annotators and as judges, would provide a more comprehensive assessment of robustness.
Finally, the kNN propagation results are influenced by the degree of overlap between the source and target questionnaires. In this paper, we evaluate kNN in a cross-standard transfer setting where the underlying control taxonomies only partially overlap, which can reduce both correctness and label diversity. An important next step is to test kNN under (i) same-distribution conditions, where new questions are sampled from the same questionnaire/standard as the labeled set, and (ii) incremental real-world conditions, where newly authored questions are added over time to an existing, partially overlapping questionnaire. These additional evaluations would clarify how much of the observed kNN degradation is due to cross-standard mismatch versus the intrinsic limitations of nearest-neighbor label propagation in sparse semantic spaces.
5 Conclusion
We presented a hybrid framework for semantic labeling in fully unlabeled compliance question repositories that combines unsupervised semantic grouping with selective use of LLMs. By inducing reusable, interpretable labels and propagating them efficiently, the approach reduces reliance on repeated LLM inference while maintaining labeling quality.
Beyond annotation efficiency, the induced labels enable more focused and adaptive generation of TPRA questionnaires. By making control domains and assessment intent explicit, the framework supports better question selection, reduces redundancy, and improves alignment with organizational requirements.
As future work, we plan to extend this framework toward automated question answering using organizational knowledge, leveraging the induced semantic labels to support consistent, reusable, and auditable responses across multiple assessments.